# Phase field modelling of the growth and detachment of bubbles in a hydrogen electrolyzer

### v1

@@ Abstract
We develop and implement numerically a phase field model for the growth and detachment of a gas bubble resting on an electrode and being filled with hydrogen produced by water electrolysis. The bubble is surrounded by a viscous liquid, has a prescribed static contact angle and is also subject to gravitational forces. We compute, as a function of the static contact angle, the time at which the bubble detaches from the substrate and what volume it has at that time. We also investigate de dependence of the detachment time on other parameters such as the applied voltage and the hydrogen ion concentration at the fluid bulk.
@@ 2.1 Fluid equations
where $\kappa(\mathbf{x},t)$ is the mean curvature at any point $\mathbf{x}$ at the surface of the bubble in contact with the liquid. We consider, as initial data, a bubble consisting of a truncated sphere with radius $R$.
The introduction of a characteristic velocity $U$ defined from the following balance between inertial and gravitational forces
@@
allows to write the system in terms of the following dimensionless numbers
$Bo=\frac{gR^{2}\rho_{1}}{\sigma},\,\,\,Re=\frac{UR\rho_{1}}{\mu},$ (5)
where $Bo$ is the Bond number and Re{Re} the Reynolds number in the liquid phase ($\rho_{1}$), as well as the density and viscosity ratios
$\gamma=\frac{\rho_{2}}{\rho_{1}},\ \chi=\frac{\eta_{2}}{\eta_{1}}.$ (6)
Hence, the velocity field satisfies
@@ 2.3 Bubble H2 influx
The pressure $p$ can be found by solving the hydrodynamic equations. Nevertheless, we can safely assume the term involving $dp/dt$ to be small compared with other terms since the pressure inside the bubble, for a bubble which grows slowly, is determined from the balance between jump in pressure across the interface and surface tension (i.e. the inverse of the bubble radius) and this changes slowly in time. Hence, the change in volume will be determined from the change in the total mass of $\text{H}_{2}$ inside the bubble.
@@ 3 Physical parameters and scales
$\mu_{1}$ $=$ $0.89\cdot 10^{-3}\ \text{Pa}\cdot\text{s},$ (39)
$\mu_{2}$ $=$ $8.874\cdot 10^{-5}\ \text{Pa}\cdot\text{s},$
@@ 4 Phase field modelling
One of the most effective methods to study multiphase flows is the so-called phase field method. The main idea is to replace sharp interfaces by “diffuse” interfaces where a phase field function $\phi(\mathbf{x},t)$ experiences sharp transitions (across the diffuse interface with a thickness of order $\varepsilon$ sufficiently small) between to limiting values (say $\phi=1$ and $\phi=-1$, for instance) corresponding to two different fluids. One of the main advantages of introducing diffuse interfaces is their ability to naturally handle topological changes in fluid domains. One needs to provide a suitable PDE to the phase field $\phi(\mathbf{x},t)$ and couple to other fluid variables in a proper way. This has been studied in many papers (cf. [7], [8], [10], [9], [11] for instance). For the phase field function, the suitable equation is a fourth order PDE called the Cahn-Hilliard equation (firstly introduced in [4], see also [3]) including a convection term with a velocity $\mathbf{v}(\mathbf{x},t)$ which is the fluid velocity:
@@
where $\mu$ is the so-called chemical potential.
In (50) $M$ is a “mobility” factor and in (51) $W(\phi)$ is a phase-field potential with the property of having two local minima at values of $\phi$ corresponding to the two phases. In particular we will take $W(\phi)=\phi^{2}(1-\phi^{2})$.
@@ 4.1 Navier-Stokes
Concerning Navier-Stokes system, we introduce the phase field as
@@
where the viscous stress tensor is given by
$\mathbf{S}=\frac{\eta\left(\phi\right)}{2}\left(\nabla\mathbf{v}+\nabla\mathbf{v}^{T}\right),$ (53)
and the material parameters $\rho\left(\phi\right)$ and $\eta\left(\phi\right)$ interpolate between the fluids densities and viscosities respectively:
$\rho\left(\phi\right)$ $=$ $\rho_{1}\phi+\rho_{2}(1-\phi),$ (54)
$\eta\left(\phi\right)$ $=$ $\eta_{1}\phi+\eta_{2}(1-\phi).$ (55)
We will also assume fluid incompressibility:
$\nabla\cdot\mathbf{v}=0,$ (56)
and for the phase field the boundary condition
@@
denoting the change of the chemical potential in the normal direction. The condition (57) is not sufficient for a fourth order equation such as (50) and we need another boundary condition. Following [9] and [10], we will impose the condition
@@
where $\sigma_{fs}^{\prime}(\phi)$ interpolates between the liquid/solid interfacial energy $\sigma_{LS}$ and the gas/solid interfacial energy $\sigma_{GS}$:
$\sigma_{fs}(\phi)=\frac{\sigma_{GS}+\sigma_{LS}}{2}+\frac{\sigma_{GS}-\sigma_{LS}}{2}\sin\left(\frac{\pi\phi}{2}\right),$ (59)
and $\sigma_{0}$ is proportional to the liquid/gas interfacial energy $\sigma_{LG}$:
$\sigma_{0}=\frac{3\sqrt{2}}{8}\sigma_{LG}.$ (60)
In this way, the condition (58) imposes, in the limit $\varepsilon\rightarrow 0$, a liquid/gas contact (or Young’s) angle $\theta_{Y}$ such that
@@ 4.2 Electrochemistry
$D_{2}(\phi)=D_{\text{H}^{+}}\frac{1+\phi}{2},\ \ \ \nabla\cdot(D_{2}(\phi)\nabla n)=0,\ \ \$(64)
@@
4.3 Bubble H2 influx
We determine the mass influx by performing a balance of mass analysis between the mass of $\text{H}_{2}$ produced at the electrode, the mass lost across the domain’s boundaries and the change of the mass of $\text{H}_{2}$ at the liquid bulk. In this way, we determine the amount of gas that enters the bubble and change its volume accordingly. We inject the corresponding amount of mass per unit time as a source inside the bubble. The flow inside the bubble diffuses the incoming gas very quickly so that we keep the gas density constant through the whole bubble except for a small neighbourhood around the source.
@@ 5 Numerical results
$J=1.24\times 10^{-10}\cdot\int_{electrode}\frac{\partial n_{\text{H}_{2}}}{\partial z}dr.$ (76)
Lastly, material properties are included for NS physics, setting the dynamic viscosity of water in $\eta_{w}=0.001\ \text{Pa}\cdot\text{s}$ and its density in $\rho_{w}=1000$ kg/m3. The viscosity and density of hydrogen are $\eta_{H_{2}}=0.00084\,\ \text{Pa}\cdot\text{s}$ and $\rho_{H_{2}}=0.089$ kg/m3 respectively, where one may notice that hydrogen viscosity is sightly greater than the real value, which was increased in order to guarantee a smooth dispersion of the injected gas and to avoid undesirable side effects, such as excessive phase field distortion produced by vortices in the bubble. The whole domain is meshed using equally-sized triangular elements of linear shape functions with a maximum element length of $0.01$ mm, which gives an average element number of 120000.
@@
The figure shows four stages in the evolution of the bubble growth, from left to right and top to bottom. At the instant $0.1$ s is shown how the angle relaxes from the initial value $\pi/3$ to the equilibrium value $\theta_{Y}$. The second plot shows the time at $t=2.06$ s, where the bubble has grown considerably. The two last images correspond to the moments right before ($t=2.10$ s) and immediately after ($t=2.11$ s) of the detachment and lift-off. Note that a small amount of gas has been left on the electrode as a remnant.
@@ 5.2 Modifying voltage and acidity
Figure 6: Lift-off times for different applied voltages. Graphs with different values of reference proton concentrations are shown in the figure, the figure shows exponential decay of the lift off time respect to the applied voltage. The effect is greater when the proton concentration at the electrode is lower.

### v3

# Phase field modelling of the growth and detachment of bubbles in a hydrogen electrolizer

@@ Abstract
We build and solve numerically a phase field model for the growth and detachment of a gas bubble resting on an electrode and being filled with hydrogen produced by water electrolysis. The bubble is surrounded by a viscous liquid, has a prescribed static contact angle and is also subject to gravitational forces. We compute, as a function of the static contact angle, the time at which the bubble detaches from the substrate and what volume it has at that time. We also investigate the dependence of the detachment time on other parameters such as the applied voltage and the hydrogen ion concentration at the fluid bulk.
@@ 1 Introduction
We remark that, in contrast to [1], the situation described here is much more complex to model. The presence of electrodes where reaction takes place and of chemical species with nontrivial boundary conditions at the surface of the bubble, introduce extra difficulties. We will implement the boundary conditions by suitably coupling the phase field with the species transport equations. The role of chemical reactions at the electrode is not only to provide the gas that fills the bubble; but to change the global geometry of the problem. As the bubble expands, including the gas/solid contact región, the área of the electrode where reaction takes place is modified and hence the production of hydrogen is influenced by the dynamics of the bubble itself. This makes the results to deviate from our previous paper due to this non trivial coupling between chemical reaction and fluid dynamics.
@@ 2.1 Fluid equations
being $\kappa(\mathbf{x},t)$ the mean curvature at any point $\mathbf{x}$ at the interface between the bubble and the surrounding liquid. We take as initial configuration for the bubble a truncated sphere with radius $R$.
We introduce a characteristic velocity $U$ defined by balancing inertial and gravitational forces:
@@
which allows us to write the system in terms of two dimensionless numbers:
$Bo=\frac{gR^{2}\rho_{1}}{\sigma},\,\,\,Re=\frac{UR\rho_{1}}{\eta_{1}},$ (5)
where $Bo$ is the Bond number and Re{Re} the Reynolds number in the liquid phase ($\rho_{1}$). Finally, defining the density and viscosity ratios
$\gamma=\frac{\rho_{2}}{\rho_{1}},\ \chi=\frac{\eta_{2}}{\eta_{1}},$ (6)
we conlcude the following equations for the pressure and the velocity field:
@@ 2.3 Bubble H2 influx
The pressure $p$ can be found by solving the hydrodynamic equations. Nevertheless, we can safely assume the term involving $dp/dt$ to be small compared with other terms since the pressure inside the bubble, for a bubble which grows slowly, is determined from the balance between jump in pressure across the interface and surface tension (i.e. the inverse of the bubble radius) and this changes slowly in time. Hence, the change in volume will be determined from the change in the total mass of $\text{H}_{2}$ inside the bubble. More precisely, we assume that the mass of $\text{H}_{2}$ produced at the electrode diffuses inside the bubble. Thus, the amount of gas across the boundaries is neglected since it is very small compared to the amount of gas produced. We will verify this by direct numerical simulations. We will inject the corresponding amount of mass per unit time as a source inside the bubble.
@@ 3 Physical parameters and scales
$\eta_{1}$ $=$ $0.89\cdot 10^{-3}\ \text{Pa}\cdot\text{s},$ (39)
$\eta_{2}$ $=$ $8.874\cdot 10^{-6}\ \text{Pa}\cdot\text{s},$
@@ 4 Phase field modelling
One of the most powerful approaches for studying multiphase flows is the phase field method. This method replaces sharp interfaces between fluids with diffuse interfaces, where a phase field function, denoted by $\phi(\mathbf{x},t)$, undergoes rapid transitions across a thin interfacial region of thickness approximately $\varepsilon$ (with $\varepsilon$ being sufficiently small). The function $\phi$ typically varies between two limiting values (e.g., $\phi=1$ and $\phi=-1$), each representing a different fluid phase.
A key advantage of using diffuse interfaces is their natural ability to accommodate topological changes in the fluid domain, such as merging or breaking of fluid regions. To model this behavior, one must define an appropriate partial differential equation (PDE) for the phase field function and couple it consistently with the other fluid variables. This topic has been extensively studied in the literature (see, for example, [7, 8, 10, 9, 11]).
The most suitable governing equation for the phase field function is a fourth-order PDE known as the Cahn–Hilliard equation, which was first introduced in [4] (see also [3]). In the context of multiphase flows, this equation is modified to include a convection term, involving the fluid velocity $\mathbf{v}(\mathbf{x},t)$, to accurately capture the transport of the phase field:
@@
where $\mu$ is the chemical potential, $M$ is a “mobility” factor $W(\phi)$ is a phase-field potential having two local minima at the values of $\phi$ that correspond to the two phases. We will take specifically,
$W(\phi)=\phi^{2}(1-\phi^{2}).$ (52)
@@ 4.1 Navier-Stokes
We introduce the phase field in the Navier-Stokes system as
@@
where the viscous stress tensor is
$\mathbf{S}=\frac{\eta\left(\phi\right)}{2}\left(\nabla\mathbf{v}+\nabla\mathbf{v}^{T}\right).$ (54)
The material parameters $\rho\left(\phi\right)$ and $\eta\left(\phi\right)$ are linear interpolation between the fluids densities and viscosities:
$\rho\left(\phi\right)$ $=$ $\rho_{1}\frac{1-\phi}{2}+\rho_{2}\frac{1+\phi}{2},$ (55)
$\eta\left(\phi\right)$ $=$ $\eta_{1}\frac{1-\phi}{2}+\eta_{2}\frac{1+\phi}{2}.$ (56)
Also fluid incompressibility will be imposed,
$\nabla\cdot\mathbf{v}=0.$ (57)
Concerning the phase field boundary conditions, we first take
@@
so that there is no flux of chemical potential through the boundary of the domain. Secondly, following [9] and [10], we additionally impose the condition
@@
where $\sigma_{fs}^{\prime}(\phi)$ interpolates between the liquid/solid interfacial energy $\sigma_{LS}$ and the gas/solid interfacial energy $\sigma_{GS}$ by means of the following formula:
$\sigma_{fs}(\phi)=\frac{\sigma_{GS}+\sigma_{LS}}{2}+\frac{\sigma_{GS}-\sigma_{LS}}{2}\sin\left(\frac{\pi\phi}{2}\right).$ (60)
$\sigma_{0}$ depends on the liquid/gas interfacial energy $\sigma_{LG}$ as:
$\sigma_{0}=\frac{3\sqrt{2}}{8}\sigma_{LG},$ (61)
so that the condition (59) imposes, in the limit $\varepsilon\rightarrow 0$, a liquid/gas contact (or Young’s) angle $\theta_{Y}$ given by
@@
We remark that the condition (59) also incorporates, in the case when the contact line is moving, and in the limit $\varepsilon\rightarrow 0$, classical models for contact line motion such as Cox-Voinov as demostrated in several studies such as [22].
@@ 4.2 Electrochemistry
$D_{2}(\phi)=D_{{+}}\frac{1+\phi}{2},\ \ \ \nabla\cdot(D_{2}(\phi)\nabla n)=0,\ \ \$(65)
@@ 5 Numerical results
$J=1.24\times 10^{-10}\int_{electrode}\frac{\partial n_{\text{H}_{2}}}{\partial z}dr.$ (77)
The flow inside the bubble diffuses the incoming gas very quickly so that we keep the gas density constant through the whole bubble except for a small neighbourhood around the source.
Lastly, material properties are included for NS physics, setting the dynamic viscosity of water in $\eta_{w}=0.001\ \text{Pa}\cdot\text{s}$ and its density in $\rho_{w}=1000$ kg/m3. The viscosity and density of hydrogen are $\eta_{H_{2}}=0.00084\,\ \text{Pa}\cdot\text{s}$ and $\rho_{H_{2}}=0.089$ kg/m3 respectively, where one may notice that hydrogen viscosity is greater than the real value, which was artificially increased in order to guarantee a smooth dispersion of the injected gas and to avoid undesirable side effects, such as excessive phase field distortion produced by vortices in the bubble. Thus the gas rapidly stabilises to a constant density state with zero velocity field. This does not change the overall dynamics of the bubble and important quantities such as the detachment time and gas volume which are the objects of our study.
The whole domain is meshed using equally-sized triangular elements of linear shape functions with a maximum element length of $0.01$ mm, which gives an average element number of 120000.
@@
The figure shows four stages in the evolution of the bubble growth, from left to right and top to bottom. At the instant $0.1$ s is shown how the angle relaxes from the initial value $\pi/3$ to the equilibrium value $\theta_{Y}$. The second plot shows the time at $t=1.5$ s, where the bubble has grown considerably. The two last images correspond to the moments right before ($t=1.72$ s) and immediately after ($t=1.73$ s) of the detachment and lift-off. Note that a small amount of gas has been left on the electrode as a remnant.
@@ 5.1 The effect of contact angle
Notice that the trend on volumen of hydrogen at detachment shown in Figure 4 deviates from the parabolic law in our previous article [1] and is more like a linear growth. There are several reasons for this. On one hand, we are dealing with a dynamical situation as opposed to the equilibrium considerations in [1]. Moreover, a balance of energies necessarily has to include, as in electrowetting phenomena, the electrostatic energy of the whole system. This is known to produce an apparent larger contact angle that is much more noticeable in drops with contact angles close to zero.
Our results are consistent with recent experimental findings (see [24] and references therein). In [24], high-speed imaging and image processing were used to analyze bubble growth and detachment in various different surfaces. The authors show that superhydrophobic surfaces significantly increase bubble volume and reduce formation frequency, while superhydrophilic and finely polished surfaces exhibit opposite trends.
@@ 5.2 Modifying voltage and acidity
Figure 6: Lift-off times for different applied voltages. Graphs with different values of reference proton concentrations are shown in the figure, the figure shows the decay of the lift off time respect to the applied voltage. The effect is greater when the proton concentration at the electrode is lower.
@@ 6 Conclusions
Acknowledgement
This work has been supported by the TED2021-131530B-I00, PID2022-139524NB-I00 and PID2023-150166NB-I00 projects, as well as URJC 2025/00014/043.